COLLISIONAL X- AND A-STATE KINETICS OF CN USING TRANSIENT SUB-DOPPLER HOLE BURN-ING

\underline{MICHAEL L. HAUSE},TREVOR J. SEARS, and GREGORY E. HALL, Chemistry Department, Brookhaven National Laboratory, Upton, New York 11973-5000.

We examine the collisional kinetics of the CN radical using transient hole-burning and saturation recovery. Narrow velocity groups of individual hyperfine levels in CN are depleted (X^{2}Σ^{+}) and excited (A^{2}Π) with a saturation laser, and probed by a counterpropagating, frequency modulated probe beam. Recovery of the unsaturated absorption is recorded following abrupt termination of an electro optically switched pulse of saturation light. Pressure-dependent recovery kinetics are measured for precursors, ethane dinitrile, NCCN, and pyruvonitrile, CH_{3}COCN, and buffer gases, helium, argon and nitrogen with rate coefficients ranging from 0.7-2.0 x 10^{−9} cm^{3} s^{−1} molec^{−1}. In the case of NCCN, recovery kinetics are for two-level saturation resonances, where the signal observed is a combination of X- and A-state kinetics. Similar rates occur for three-level crossover resonances, which can be chosen to probe selectively the hole-filling in the X state or the decay of velocity-selected A state radicals. However in the case of CH_{3}COCN, which has a dipole moment of 3.45 D, the X-state kinetics are faster than the A-state due to an efficient dipole-dipole rotational energy transfer mechanism as the X-state dipole moment is 1.5 D and the A-state dipole moment is 0.06 D. The observed recovery rates are 2-3 times faster than the estimated rotationally inelastic contribution and are a combination of inelastic and velocity-changing elastic collisions.

Acknowledgement: This work was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy.